Avending machine contains nickels and dimes worth $14.50. There are 95 more nickels than dimes. Hom many nickels and how many dimes are there?
Answer by ewatrrr(24785)


Hi
There are 95 more nickels than dimes.
Let x and (x+95)represent the number of dimes and nickels respectively
Question states*** CENTS makes sense
10x + 5(x+95) = 1450
15x = 975
x = 65, the number of dimes. Ther is 160 Nickels (65+95)
CHECKING our Answer***
65*10 + 160*5 = 650 + 800 = 1450 or $14.50